e table 20) Find 2o)
Answer #1

1. a) i) P( \chi^2_{28 } >= X) =0.005

df = 28

At \alpha = 0.005 and df = 28, from chi-square table, X= 50.993

or using excel function = CHISQ.INV.RT(0.005,28) = 50.993

---------------------

a) ii) P( \chi^2_{20} <= X) =0.95

= 1-P( \chi^2_{20} > X) = 0.95

= P( \chi^2_{20} > X) = 0.05

df = 20

At \alpha = 0.05 and df = 20, from chi-square table, X= 31.410

or using excel function = CHISQ.INV.RT(0.05,20) = 31.410

---------------------

b) i) P(\chi^2_{20} <= 44)

= 1-P(\chi^2_{20} > 44)

Using excel function:

=1-CHISQ.DIST.RT(44,20) = 0.9985

-----------------

ii) P(\chi^2_{33} <= 41)

= 1-P(\chi^2_{33} > 41)

Using excel function:

=1-CHISQ.DIST.RT(41,33) = 0.8402

------------------------

If you have any doubt ask me in comments.